1. If the discounted amount is ₱8500 while the discount rate is 22% then what is the future amount of the money?




2. Compute the annuity of the money which has future worth of ₱50 000 in 5 years if it compounded monthly at the rate of 11%?






3. What is the equivalent effective rate of interest compounded quarterly if the nominal rate is 14.5%?


1
Expert's answer
2022-03-24T16:09:00-0400
  1. discounted amount = 8500

discount rate = 22%

future value = 100%

"100-22=78"

78% = 8500

100% = x

"78% * x=" "*x= 8500*100"

"78x=850000"


"\u200b\n \\frac{78x}{78} = \n\n\u200b\n \\frac{850000}{78}"


"x=10897.4359"

The future value is approximately P10897.


2.Future value (FV) = 50000

time n= "5*12=60"

interest rate = 11%

monthly interest rate = "11\/100\u00f712""=" 0.009166

present value (PV)

"PV = FV[\frac{1}{(1+r)^n}" ]

"(1+r)^n= (1+0.009166)" 60 =1.728916


"PV= 50000[\frac{1}{1.728916}]"


= 28919.8595

The present value is approximately 28919.86


3.

Nominal rate (r)= 14.5%

Compounding (m)= 4/yr

effective rate interest (i)

"i= (1+\frac{r}{m}" ) m -1


"i=(1+\frac{0.145}{4}")4 -1


"i=0.15307664"

I= "i*100"

I"= 0.15307664*100"

= 15.307664

The effective interest rate is approximately 15.3077%